Fermi National Laboratory
 
Operation Summary for December 21, 2011 - December 28, 2011
 

NuMI reported receiving 8.81E18 protons on target for the period from 12/19/11 to 12/26/11.

Wednesday December 21

The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, and with the MiniBooNE and NuMI taking beam.

During the day shift, Operators reported some downtime due to Linac personnel changing out an analog to digital chassis and due to Water Group personnel switching MI RF cavity LCW pumps. There was also some downtime when Operations rebooted MECAR at the request of a MI person. Safety System personnel gave Operations a Special Operating Note for Muon Rings high intensity studies.

During the evening shift, a Linac RF station (LRF3) tripped off several times on a permanent inhibit status; it reset. The NuMI beam permit dropped when a NuMI BPM (NBPM65) stopped polling; it reset. The I- Source started to wander; Operators made adjustments.

Thursday December 22

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, a Muon Rings person began a high intensity study at 5:02 AM. Operators rebooted NBPM65.

During the day shift, Muon Rings high intensity studies finished at 9:30 AM. Operators tuned Booster to improve future high intensity Muon Rings studies. Booster and Recycler studies followed the Muon Rings studies. At 1:27 PM, a Booster kicker (B:MKS08) tripped off; Booster personnel repaired a cable fault and brought the station back online.

During the evening shift, the Meson FTBF T-1019 experiment began taking beam. Operators adjusted the LRF3 driver screen voltage. An I- Source vacuum pump tripped off; it reset.

Friday December 23

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, a Booster RF station (BRF5) tripped off. Operators started to investigate when the facility system issued a smoke alarm. Operators prepared the area for an access; Firefighters entered the enclosure and found that the smoke had came from a component that had suffered from a voltage surge. Later, Booster personnel reported finding a transformer that had a coil short, which caused the fault in the tunnel; they began repairs.

During the day shift, Booster beam went off at the end of the midnight shift for repairs to BRF5. Booster resumed sending beam to NuMI and MiniBooNE at 1:50 PM. Operations established beam to T-1019 at 2:50 PM. Operators reported problems with BRF4. Booster personnel replaced the station’s modulator, which took down beam as the station was tied in and out. Water Group personnel conducted Tevatron pond water pump tests.

During the evening shift, Operators replaced a NuMI CAMAC crate power supply. Operations disabled T-1019 beam at 10:02 PM.

Saturday December 24

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, Operations monitored the accelerator complex.

During the day shift, Operations established T-1019 beam at noon. Operators reported several Linac and Booster RF trips that caused short downtimes during the shift.

During the evening shift, Operations disabled T-1019 beam at 10:01 PM.

Sunday December 25

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, Operations monitored the accelerator complex.

During the day shift, Operations established T-1019 beam. BRF17 tripped off throughout the day due to 'timer delay active' and 'deck power off' indications; these trips took down the West Anode Power Supply. Operators bypassed the station and reconfigured the RF system..

During the evening shift, around 5:45 PM, the I- Source starting to crash; Operators responded by increasing its pressure and the Source recovered. Operators tuned the Source’s magnet current, the 750 KeV line, and Linac. Operations disabled T-1019 beam at 10:01 PM.

Monday December 26

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, Operations monitored the accelerator complex.

During the day shift, Operators reported a short downtime to bring BRF17 back on line. Operations disabled T-1019 beam at 3:30 PM.

During the evening shift, Operators reported some minor I- Source problems. Operators also rebooted NBPM65 several times. Operators tuned the Booster and MI accelerators.

Tuesday December 27

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, Controls personnel reported that they were replacing a Switchyard (SY) CAMAC crate power supply.

During the day shift, Operations reported that from 9:39 AM to 12:32 PM, Booster Personnel conducted aperture scan studies. Operations established beam to T-1019 at 11:06 AM. At 3:30 PM, Recycler personnel began a study.

During the evening shift, the Recycler study ended at 5:08 PM. Operators reported problems with MIRF11; they investigated and bypassed the station. Operations disabled T-1019 beam around 10 PM.

Wednesday December 28

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, Operations reported that a scheduled Booster beam intensity reduction study started at 4 AM and finished at 4:50 AM. Operators reported that they had to reconfigure MIRF15 to compensate for MIRF11 and MIRF17 being off. A RF person brought both stations back on, however, not to full gradient. He was still investigating at the end of shift.

Operations Coordinator (OpsCo)

The OpsCo approved of the plans by Booster and MI personnel to conduct studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/28/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory